Speccy:

Welp, I'm dead! 7/6d + 4/3 = -1/3

annacam:

𝑑 = − 1 0 7 d= -10/7

Sailor:

Im pretty sure you would multiply the fractions first-

Speccy:

I have never seen this before. It's gong to be a while before I understand.

Sock:

why r there so many people on this post? lol

Speccy:

21/24

Speccy:

What am I doing first yall?

Astrid1:

Alright so, this is our equation. \[\frac{ 7 }{ 6}d+\frac{ 4 }{ 3 }=-\frac{ 1 }{ 3 }\] Subtract 4/3 from each side. \[\frac{ 7 }{ 6 }d+\frac{ 4 }{ 3 }-\frac{ 4 }{ 3}=-\frac{ 1 }{ 3 }-\frac{ 4 }{ 3 }\] Simplify: \[\frac{ 7 }{ 6 }d=-\frac{ 5 }{ 3 }\] Multiply each side by 6 \[6\times \frac{ 7 }{ 6 }d=6~-(\frac{ 5 }{ 3 })\] ^^ parenthesis goes in front of the 5 & the 3. Just wont let me add it there, I apolgize. Simplify 7d=-10 Divide both sides by 7 \[\frac{ 7d }{ 7 }=\frac{ -10 }{ 7 }\] ^negative goes in front of 10 and 7 again ;c Simplify: d= -10/7

Astrid1:

@speccy

Speccy:

Oh! Thanks Swol.